The measure of an angle is eight times the measure of a complementary angle. What is the measure of each angle?

Answer by ewatrrr(24785)


Hi
Let x represent the larger angle
x = 8(90°-x)
x = 720 - 8x
9x = 720°
x = 80° and it's complement is 10°
and...checking
80° = 8*10°